OpenStudy (anonymous):

one leg of a right triangle measures 8√ 6 units and the other leg measures 8√ 2 how long is the hypotenuse? A.8 B.16 C.16√ 2 D.16√ 3

OpenStudy (anonymous):

use pythagores theorem

OpenStudy (anonymous):

Using the Pythagorean theorem, we know that c^2 = a^2 + b^2. So pluging in the numbers, we have : c^2 = (8sqrt6)^2 + (8sqrt2)^2 so, have you seen how to multiply radicals?

OpenStudy (anonymous):